OpenStudy (anonymous):

write an equation in the form y=mx+b for the line which passes through (1,-10) and (-7,-2)

pooja195 (pooja195):

First we need to find the slope any ideas on how to find it?

OpenStudy (anonymous):

ok know how to do that give me a min

OpenStudy (anonymous):

slopes -1

pooja195 (pooja195):

ok next are you fimilar with the point slope formula?

OpenStudy (anonymous):

not really can you remind me

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\] m=slope int his case -1 next we select any random post lets go with (1,-10) so 1 = x1 -10=y1 now you just plug it into the formula

pooja195 (pooja195):

Are you ok with that? Do you know how to do it?

pooja195 (pooja195):

stuck? :/

OpenStudy (anonymous):

a little

pooja195 (pooja195):

\[\huge~y--10=-1(x-1)\]

pooja195 (pooja195):

therefore we have \[\huge~y+10=-1(x-1)\] now distribute \[\huge~-1(x-1)\]

OpenStudy (anonymous):

8

pooja195 (pooja195):

huh?

OpenStudy (anonymous):

hold on messed up bad not smart at algebra at all

pooja195 (pooja195):

need help? :)

OpenStudy (anonymous):

yes i need all the help i can

pooja195 (pooja195):

therefore we have \[\huge~y+10=-1(x-1)\] now distribute \[\huge~-1(x-1)\]

pooja195 (pooja195):

distribute :)

pooja195 (pooja195):

-1 * x = ? -1 * -1 =?

OpenStudy (anonymous):

any thing that x right

pooja195 (pooja195):

-1x :) -1 * -1= ?

OpenStudy (anonymous):

1

pooja195 (pooja195):

good so we have \[\huge~y+10=-1x+1\] we are trying to isolate y so what would the next step be?

OpenStudy (anonymous):

so what your saying is get rid of 10 or change y

pooja195 (pooja195):

get rid of the 10

OpenStudy (anonymous):

do i subtract 1 from 10

pooja195 (pooja195):

yes so 1-10= ?

OpenStudy (anonymous):

-9

pooja195 (pooja195):

good so we get \[\huge~y=-1x-9\] thats all :)

OpenStudy (anonymous):

thank you thank you thank you :)

pooja195 (pooja195):

:)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions